OpenStudy (anonymous):

Quick help gets medal! How many solutions are there to the following system of equations? 2x + 5y = –11 10x + 25y = –55 A. 2 B. 1 C. infinitely many D. 0

OpenStudy (kamille):

You should solve the system of equation and check how many solutions it has. Do you know how to do it?

OpenStudy (anonymous):

NO help please

OpenStudy (kamille):

NO, do it yourself.

OpenStudy (anonymous):

@Hero

hero (hero):

When you divide the second equation by 5, you get 2x + 5y = -11 Which is the same as the first equation. Do you know what that means?

OpenStudy (anonymous):

sorry caps was on @Kamille

OpenStudy (anonymous):

They're equal

hero (hero):

So if both equations are equal, how many solutions are there?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?

hero (hero):

I don't want you to guess. Please try out different values. Try (1, 2) and (2,1) and see what happens.

OpenStudy (anonymous):

I don't get it?

OpenStudy (anonymous):

So im supossed to try more numbers?

hero (hero):

I'm only going to tell you this once....don't forget it... If both equations are the same, then there are infinite solutions. Never forget this.

OpenStudy (anonymous):

ok thanks ill never forget.

hero (hero):

When in doubt, try graphing both lines.

OpenStudy (anonymous):

Ok but I have one question?

OpenStudy (anonymous):

how did you get the 5 to divide the 2nd equation.

hero (hero):

Basically, the first thing I do before anything anything else is to see if I can reduce either equation by factoring. Notice that 5 factors out of the second equation since the coefficient of each term is divisible by 5: 10x + 25y = –55 5(2x + 5y) =5(-11) Then after dividing both sides by 5 you get: 2x + 5y = -11